NAOJ GW Elog Logbook 3.2

Yuhang and Michael
We measured again the RIN of the new Mephisto laser. The laser was sent to a tilted ND filter (prevents back reflection) -> 75mm focal length lens -> PD. The measurements were taken with the spectrum analyser taken out of the clean room.
The relative intensity noise and PD dark noise are shown in figures 1 and 2 (low freq span, high freq span). The units differ between the curves, though the RIN isn't recalibrated by much - the intensity noise was measured from the PD and then divided by the DC value of 0.89 V. For comparison, the KAGRA HP laser RIN shown in Aso-san's LVK presentation is given in figure 3. The new Mephisto laser at TAMA seems to have better RIN than the "Current" KAGRA laser, but not quite as good as the "New" laser.

Today I checked the 2 PBS available in PCI.
There is one with label (1064 nm PBS) so I decided to use this one.
I moved it to the mount with the good height but I'll need to print a new label for it.
I measured its transmission to be 166.0 uW with 166.4 uW of incident power.
Actually, it seems that the beam is too large for the small power-meter head as the power is supposed to be 30 mW there...
Anyway I started to tweak the alignment by checking the AC value of the surface reference sample.
Up to now I did not manage to recover proper alignment and it seems that the shape of the AC signal during Z scan is a bit strange (kind of similar to bulk sample...)
Hopefully I'll have a bit more time tomorrow to finish this alignment.
Note that to save time the pump laser is kept on with low power (30 mW)

I attached the measurement results which gives absorption = 267+/-59 ppm/cm with P_in = 3.745 W and P_t = 3.198W.

Yuhang and Michael
We did a quick measurement of the new incoupling mirror using the OPO replacement setup in ATC. The incoupling mirror was placed with the convex side facing forward in the mirror mount. The beam was reflected at a low angle (approximately 10 degrees), which is within the specified incident angle tolerance. Reflection and transmission were measured on both faces of the incoupler.
Using thorlabs power meter, we obtained:
Forward mount
Incident = 6.58 mW
Reflected = 6.07 mW (92.2%)
Transmitted = 0.537 mW (8.1%)
Reverse Mount
Incident = 6.64 mW
Reflected = 6.04 mW (91.0%)
Transmitted = 0.535 mW (8.1%)
The power meter is imprecise when being intermittently moved around between measurements - certain beam spot positions around the edge of the power meter seem to record more power (also noted during a previous measurement of OPO nonlinear gain in TAMA)
Very rough measurement obtained by focusing ceiling lights onto the table indicates the focal length to be approximately 15cm.

Today I installed the SHINKOSHA evaluation plate #7 on the translation stage so that the top side of the ingot is facing the laser sources with marking on the right side of the sample looking from the laser source side.
I moved the IU to 62.6 mm to compensate for the thickness.
I checked the X and Y centering using drop of the DC values to ~ 8mV which gave : X_center = 397.466 mm and Y_center = 120.629 mm.
I increased the pump power to ~3W and checked the Z centering at that position from the phase = 0 deg and got Z_center = 53.2 mm.
I also checked the top/ bottom and left/right Z centering over a 120 mm diameter area and found a tilt of this sample in both these directions to be ~ 0.8 mm.
Matteo confirmed that it was small enough to start absorption measurement.
The XY absorption measurement started over a circular area with diameter 120 mm and step size 250 um with P_t = 3.198 W and P_in = 3.745 W

Mainly two mechanisms cause zero baseline drift (ZBD): the birefringence of electro-optic crystal and the etalon effect formed by two parallel end facets of the EOM crystal [Z. Li, et al., Optics Letters, 41, 14, 2016]. The reduction of birefringence effect can be done by controlling modulation voltage or the crystal's temperature [K. Kokeyama, et al., J. Opt. Soc. Am. A 31, 81 (2014)].
To monitor ZBD, we can lock cavity on anti-resonance. In this situation, the PDH signal is very much insensitive to phase change. Therefore, the signal magnitude change comes mainly from sideband amplitude change, usually addressed as residual amplitude modulation (RAM). This amount of RAM exists all the time, including the situation when cavity is on resonance. The control loop makes error signal to be zero with such RAM present, which makes the cavity not exactly on resonance. Or we can say RAM introducing detuning for cavity. This is the effect we are interested in. Notably, this effect may appear also for the CCFC control loop, which should be investigated.
We used the method described in the last paragraph to monitor ZBD. Since this effect exists when cavity is on resonance, we use the slope of PDH signal at resonance to calibrate ZBD. In this way, we have the detuning influence cause by ZBD for detuning change. The attached figure (is not yet calibrated), after calibration, shows a ~4Hz detuning change around zero.

Measurement finished and reported in the 3 attached figures.
I finally choose to use this colormap for all plots as it is (I think) nice looking, color-blind and printing friendly.
Mean absorption seems to be 18+/-5 ppm/cm.
The sample was oriented to have the kanji for 'up' at the top.

Two HEPA capsules are installed for the air gun of PCI cleanroom. The maximum air pressure was specified to be 3.4 bar (figure 1). So I adjusted the pressure of the compressed air as attached figure2.
History: Marc found that there is air leakage between the connection part of capsule and tube. The connection was done by only fixing capsule and tube with steel belt. Aso-san put some white sealing tapes in this connection part. The use of sealing tape solves the problem of air leakage.

Yesterday I started the measurement of the tama size sample T1.
Thanks to the help of Simon and Matteo I could identify the good sample.
Indeed there are now 2 tama size sample in PCI clean room but one seems to be different (maybe FC mirror?)
The one I installed is the one on the left of entry 301.
However, it seems that this sample creates at least 2 beams in reflection that hit directly the last prism on the probe path before the sample.
In turn, multiple reflections are sent a bit everywhere (but not towards the photo detector).
Using 50 mW of incident pump power I could estimate the power lost in these reflections to be around 10% of the incident power.
Because the sizes of these reflections are quite large, we started the measurement anyway.
These measurements will finish this evening and I will turn off the pump laser at that time.
Measurement finished and reported in the 3 attached figures.
I finally choose to use this colormap for all plots as it is (I think) nice looking, color-blind and printing friendly.
Mean absorption seems to be 18+/-5 ppm/cm.
The sample was oriented to have the kanji for 'up' at the top.

The 3 measurements (XY, XZ and YZ) have been finished and are reported on the attached figures.
Note that for the XZ and YZ measurements the mean and standard deviation are computed only inside the sample while the contour plot and histogram show the entire measurement.
The results is compatible with previous measurements and gives absorption ~ 60 +/- 5 ppm/cm without peculiar feature on XY plane.
I like a lot the plots in gray scale but I also did the same with colored scale.
I'll upload soon a notebook to do these plots.
The mirror is oriented to have a black dot on the top.
EDIT : changed number of bins in histogram

Yuhang, Michael
We measured the relative intensity noise of the new Mephisto 2W NdYAG laser to be used at TAMA. The laser was sent into a 75mm lens, then a 50:50 beamsplitter going to a photodetector and power meter sensor. The distance between the laser and detectors is fairly short due to the initial divergence of the laser beam. With the photodetector we took the following spectra, while the power meter was used to measure long term power stability. Some ND filters were placed between the laser and beam splitter to counteract saturation of the PD and prevent back reflection into the laser.
First PD: Thorlabs PDA36A-EC
Second PD: Qubic PD-AC100-Si 1-100 MHz
Figure 1 shows the low frequency laser RIN at approximately the minimum current required for the laser to start emitting (about 0.8 A), compared to the PD dark noise. The measurement was passed through the Stanford pre-amp at DC coupling and 100 gain.
Figure 2 shows the RIN at different laser current (power). The DC level on the detector changes on the highest two measurements due to the PD saturating. The gain on the pre-amp was reduced by a factor of half for the highest two current measurements. Due to the PD reading being close to the noise floor, this may be responsible for the jump between the levels of the measurements between different pre-amp gain values.
Figure 3 shows the effect of activating the noise eater. In our first measurement its effect only became noticeable at laser current > 1.55 A.
Figure 4 shows the predicted frequency dependence of RIN from the Coherent Mephisto manual. Ideally the noise eater suppresses the large peak and reduces the 10k-100kHz noise floor by ~15 dB. We don't see much if any of the latter effect though. It seems like the low frequency signal is too close to the detector noise level. The levels are a bit different from the measured results - I didn't apply normalization to the measurements from the spectrum analyzer readings, but the relative values for noise eater on/off are what is important.
To follow up, we tried changing the preamp gain settings and changing the spectrum analyser coupling from DC to AC to see if we could get better SNR. The Stanford pre-amp was close to saturating though.
Figure 5 shows the result of increasing the PD gain and optimising the pre-amp settings. We can get a bit more clearance from the PD dark noise versus figure 1, but still less than desireable.
We then tried to do another measurement using the second PD, but it seemed like it had more dark noise than the Thorlabs one just from inspection on the oscilloscope. The Qubic PD is AC only, but we couldn't see anything on the oscilloscope when applying 1-10 kHz modulation, just ~ 5 mV rms noise.

Today I checked the bulk reference and got : R_bulk = 0.7527 cm/W (note that in this case I moved the IU by 0.32 mm)
With laser diode current of 6A I set up the HWP so to have P_in = 7.072W.
Using the DC values drop, the Kasi sample center positions are : X_center = 329.62 mm and Y_center = 122.81 mm.
The Zaber limits have been changed accordingly with some margin to check the incident power from remote.
I did a long Z scan and used the phase transition to get the center Z position at Z_center = 64.1 mm.
I realigned the DC at this position and started a circular map in the XY plane with parameters :
X = X_center, Y=Y_center, Z =Z_center, 20 mm radius, 0.1 mm step size, 500 ms waiting time, median/average filter order 5 and sensitivity = 5mV (ie 5 nA).
Preliminary absorption is : 75 ppm/cm (coherent with previous measurement)
with :
AC =4.6e-4 V
DC = 4.144 V
P_in = 7.074 W
P_t = 6.096 W
T_sample = P_t/P_in;
mat_correction=3.34;
All measurements should be finished tomorrow.

Katsuki, Marc
We performed several scans of the the z positions of the translation stage or the IU with the surface reference sample.
We plan to check if the shape of these signals could be used to perform more efficiently the alignment.
We tried to place back the PBS for the birefringence measurement but it was quite lossy (~20%).
Actually the one we used does not have name but the holder height seems to match. So we installed it as in entry 1496.
I'm planning to buy a new PBS as I could not find other.
The R_surf is now 16.79 /W at z_translationStage = 40.5 mm and z_IU = 66 mm.
Tomorrow we'll start the absorption measurement of the korean sample.

Michael and Yuhang
We received a new Mephisto laser this Monday. Michael and I tested lasing current threshold, noise eater engagement current threshold. We measured RIN with different frequency band, current, noise eater on/off, and power stability. In this entry, we report new Mephisto power stability.
We used a power meter to monitor the power stability of new Mephisto. The total monitoring time is 17.5 hours. The power change is attached in the figure. A peak to peak 1.4% power stability was found with irregular power change found at the end of measurement. This shows that laser power doesn't become really stable after long time operation.

The IR detuning map in elog2615 was concerned that it could be related with alignment control. Therefore, I try to clarify this concern here and do some related test.
How is the AA loop?
As far as I found, the AA loop bring GR and IR transmission to top of TEM00. If I introduce pertubation to AA loop, GR power goes down, which means AA loop works well.
Why IR detuning map should not be related to AA?
1. The IR detuning map in elog2615 should not be related to AA loop because the AA loop works to have a totally linear response as elog2650. But IR detuning map has a flat region.
2. I took IR detuning spectrum with AA loop opened. Meanwhile, the AA error signals are put around zero by hand. The beam hitting position was chosen as elog2573 to old and new spot. Then I got two spectrums are the attached figure. The spectrum shows result as found in elog2573. Since AA loop is open, there should be not effect from AA loop.

It is found, as attached figure 2, that PR and BS mirror had a sudden position change on last Thursday (20210812). This makes the filter cavity alignment totally lost. Since the movement of PR mirror is so large that the PR mirror picomotor needed to be used to recover alignment.
The mirror movement has coincidence with an earthquake (figure 1). But it is the first time I notice that mainly PR pitch mirror is moved. BS is moved a bit as well. But no obvious mirror movement is found for input and end.

The tire of the bicycle in the south arm was broken. However, the one in the west arm was OK. So I exchanged the place of them. In the future, we can use a good bicycle (blue color as attached picture) to go to the end room of south arm.

I measured CCFC error signal for 3 hours with 30 minutes interval (Fig. 1). Fig. 2 shows the CCFC error signals around 0 crossing point. The 0 crossing point of the fitting result (dashed curves) changed by (1.04-0.97)*54 = 3.8 Hz in 3 hours.
I also measured the nonlinear gain and shot noise before/after the whole measurement.
The nonlinear gain changed from 4.7 to 4.3 in 3 hours. This nonlinear gain change causes the CCFC amplitude change. The normalized CCFC amplitude can be written as follows.
Normalized CCFC amplitude = x/(1-x^2)^2 = (1-1/sqrt(g))/(2/sqrt(g)-1/g)^2
Figure 3 shows the normalized CCFC amplitude as a function of nonlinear gain. When the nonlinear gain changes from 4.7 to 4.3, the normalized CCFC amplitude changes from 1.07 to 0.97 by a factor of 0.9. In fact, the CCFC amplitude changed from 132mVpp to 118mVpp by a factor of 0.9 in 3 hours.

I measured CCFC error signal with different length LEMO cables for CCFC LO (attached figure). The CCFC amplitude was 132mVpp. The red and green curves in the figure represent the CCFC error signal with red and brown+green LEMO cables for CCFC LO, respectively. As you can see, the CCFC error signal with brown+green LEMO cable is close to I phase. So I will use the brown+green cable for CCFC LO.

I measured CCFC FDS with fixed homodyne angle for 3 hours with 30 minutes interval (figure 1). The FC was unlocked between each FDS measurement. According to the least square fitting, the detuning changed by 7 Hz in 3 hours even with fixed homodyne angle.
Before each FDS measurement, I optimized p pol PLL frequency to have maximum BAB transmission with 20mW green. The nonlinear gain change was 4.4-4.7 in 3 hours, which corresponds to generated squeezing of 10.1-10.5 dB. Since I optimized p pol PLL every time I measured the nonlinear gain, this nonlinear gain change is the real nonlinear gain change, not the detuning change of BAB.
I also measured shot noise before/after all the FDS measurement (figure 2). The shot noise changed by 0.15dB in 3 hours.
sqz_dB = 10.5; % generated squeezing (dB)
L_rt = 120e-6; % FC losses
L = 0.49; % propagation losses
A0 = 0.06; % Squeezer/filter cavity mode mismatch
C0 = 0.02; % Squeezer/local oscillator mode mismatch
ERR_L = 1.5e-12; % Lock accuracy (m)
ERR_csi = 30e-3; % Phase noise (rad)